OpenStudy (anonymous):

What is the solution to the system of equations? y = x + 3 x = –2

OpenStudy (vhtran):

Instead of having "x" in the equation, replace it with -2

OpenStudy (vhtran):

if you do that you get y=-2+3

OpenStudy (shadowlegendx):

y = x + 3 x = -2 As @vhtran said, replace the x with -2 \[y = -2 + 3\] What is -2 + 3? That is your answer :0